OpenStudy (anonymous):

how many solutions does 2=2 have?

OpenStudy (danjs):

infinity

OpenStudy (danjs):

k*2=k*2 for k = anything

OpenStudy (anonymous):

any number is set equal to the same number...yeah..there should be infinite solutions